OVHCloud Web Hosting Status

Current status
Legend
  • Operational
  • Degraded performance
  • Partial Outage
  • Major Outage
  • Under maintenance
FS#8972 — Optical loop Lille
Scheduled Maintenance Report for Web Cloud
Completed
Tonight around 1am we will upgrade the optic equipment version in Northern France.

This upgrade aims to support 10G lambdas, no downtime planned. This operation is transparent.

Update(s):

Date: 2013-07-26 02:15:02 UTC
POP2 OK.

Date: 2013-07-26 02:14:56 UTC
hau ok, doing pop2.

Date: 2013-07-26 02:14:44 UTC
rsa and stm ok, doing hau and can..

Date: 2013-07-26 02:14:24 UTC
ron and tou ok, doing rsa and stm..

Date: 2013-07-26 02:14:05 UTC
was and san ok, doing ron and tou..

Date: 2013-07-26 02:13:40 UTC
tri and hel ok, doing was and san..

Date: 2013-07-26 02:12:40 UTC
lzi and asc ok, doing tri and hel..

Date: 2013-07-26 02:12:27 UTC
pil and lqn ok, doing lzi and asc..

Date: 2013-07-26 02:12:17 UTC
ble59 ok, doing pil59 and lqn59..

Date: 2013-07-26 02:12:05 UTC
boi59 ok, doing ble59..

Date: 2013-07-26 02:11:59 UTC
fac59 ok, doing boi59..

Date: 2013-07-26 02:11:17 UTC
sec59 ok, doing fac59..

Date: 2013-07-26 02:11:08 UTC
lib59 ok, doing sec59..

Date: 2013-07-26 02:11:01 UTC
pop-1 ok, doing lib59..

Date: 2013-07-26 02:10:51 UTC
upgrading service cards of pop-1.

Date: 2013-07-19 02:10:32 UTC
We delayed the intervention on Lille's optical loop to Monday night.

Date: 2013-07-17 23:30:19 UTC
Thereafter the issue on this task http://status.ovh.net/?do=details&id=5050, we are postoponing the intervention on the optical loop Lille to tomorrow Friday the 19th at 01:00.


Date: 2013-07-17 23:27:27 UTC
upgrading pop-1 modules.

Date: 2013-07-17 23:27:10 UTC
rebooting pop-1

Date: 2013-07-17 23:27:03 UTC
we will start with pop-1
Posted Jul 17, 2013 - 10:00 UTC